Project

General

Profile

Kaon LT Meetings » mtg_22apr19.txt

Garth Huber, 05/01/2022 11:05 PM

 
1

    
2
		Apr 19/22 PionLT/KaonLT Analysis Meeting Notes
3
                ----------------------------------------------
4

    
5
Participants:
6
 - Tanja Horn (arrived late), Richard Trotta, CUA
7
 - Peter Bosted, Dave Gaskell, JLab
8
 - Jacob Murphy, Julie Roche, Ohio
9
 - Pete Markowitz, FIU
10
 - Nathan Heinrich, Garth Huber, Muhammad Junaid, Stephen Kay, Vijay Kumar,
11
   Love Preet, Ali Usman, Regina
12

    
13
Please remember to post your slides at:
14
https://redmine.jlab.org/projects/kltexp/wiki/Kaon_LT_Meetings
15

    
16
Richard
17
-------
18
- 10.6 GeV Heep-SINGLES analysis and SIMC comparison
19
  - adjusting offsets by hand, trying to determine a consistent SHMS offset for
20
    ssdelta
21
    
22
- *Suggestions*
23
  PB: has determined offsets for all of our KaonLT data
24
      - shifts delta in both hcana and SIMC, where offsets are stored in a
25
        master file
26
      - NOTE: finding a good set of offsets is sometimes tricky, as there can
27
        be multiple sets that give similar results
28
      - it would be good for Richard to compare his eventual offsets with PB's
29
        to see if they are different, or not
30
  DG: will send a copy of Henk's Heep Check Offsets Program
31

    
32
- otherwise continuing to deal with UTIL_KAONLT repo issues
33

    
34
- *Goal for Next Meeting*
35
  - a set of 10.6 GeV offsets
36

    
37
Ali
38
---
39
- fixing Heep Singles settings in standard.kinematics
40
  - results will be pushed soon, about 2/3 done
41
  - some 10.6 GeV angle TV settings were not recorded.  Will have to get the
42
    photos from raw data
43

    
44
- *Goal for Next Meeting*
45
  - either have all of the physics settings fixed in std.kin OR look at Heep
46
    8.2 data/SIMC comparison, depending on how fast things go
47
  
48
Vijay
49
-----
50
- PionLT low Q2 preliminary data/SIMC comparisons
51
  - 2.7 GeV done
52
  - 3.6 GeV: one setting done
53
  - 4.5 GeV; not yet started
54

    
55
  - still using OLD offsets
56
  - only needs to rescale SIMC*0.8, while low Q2 KaonLT needed 7x
57
  - comparisons generally look good
58

    
59
- *Suggestions*
60
  PB: xptar, yptar data/SIMC distribution edges look different
61
      *it is very important to have data and SIMC edges match*
62
    - had a similar problem in his analysis
63
    - shms_hut, hms_hut in SIMC has a flat for 3/4 scintillator trigger
64
      simulation
65
    - seems that the SIMC routine has a tighter scintillator grid than the
66
      actual detectors, for example the particles also give a Cherenkov signal
67
      in the hodo lightguides that is excluded in SIMC
68
    - to get around this problem, he turned off the flag in SIMC to ignore
69
      hardware cuts, then applied cuts later in both hcana and SIMC
70
    - the cuts that were applied were:
71
      - projection to S1, S2, aero, shower (fiducial cuts)
72
      - PB will send a copy of the fiducial cuts he used 
73

    
74
   DG: does not like this approach
75
    - would prefer to fix any problems in SIMC, rather than turn off this flag
76
    - we need to be sure that the hodoscope size and positions in SIMC
77
      are correct
78
    - TH agrees that this is the previous procedure, greatly preferred
79

    
80
- *Goal for Next meeting*
81
  - coplete rest of of PionLT low Q2 settings, then will go back to low Q2 Heep
82
    settings, and start to fix problems indentified in the preliminary analysis
83
    
84
Jacob
85
-----
86
- new SHMS hodo calibrations for Run 12055
87
  - major fix: hard cut on paddle constraints adjusted by Dave Mack
88
  - new velocity constraints by Carlos Yero
89
  - time walk TDC threshold parameter raised 120 -> 1200
90
    - p.hodo.beta distribution much improved (plot by CY)
91
    - trying to replicate CY's improvement
92
  - 2D plot p.gtr.beta:p.gtr.dp
93
    - weird blob structure and wiggles gone
94

    
95
- *Goal for Next meeting*
96
  1) fits for all different fall 2021 run settings
97
    - already looked at defocused Run 11799
98
  2) look at HMS hodo calib to see if any improvements result
99
    - HMS previously did not exhibit the same issues as the SHMS hodo calib
100

    
101
Junaid
102
------
103
- just got his JLab account back
104

    
105
- *Goal for Next meeting*
106
  - drift chamber and calorimeter calibrations
107

    
108
Nathan
109
------
110
- no report.  working on PionLT run plan and QFT class
111

    
112

    
113
Stephen
114
-------
115
- no report.  will look at online software soon
116

    
117
Dave G
118
------
119
- reported that no one (MJ, electronics group) is enthusiastic to do the FADC
120
  reprgramming, so it will be the same as the last run (i.e. we will not get
121
  the conventional data when there is an error, only Mode-10, so it will not be
122
  possible to directly compare them)
123

    
124
- Brad S is switching positions from A/C to Computer Center at end of month
125
  - will still be available for questions should we need him
126

    
127
Garth
128
-----
129
- will draft a call for shift workers
130

    
131
Next Meeting
132
------------
133
- Tues May 3 at 14:00 Eastern
(41-41/412)